# Vietnamese Poem Generation & The Prospect Of Cross-Language Poem-To-Poem Translation

### 03 Jan 2024 v2

@@ Abstract.
Poetry generation has been a challenging task in the field of Natural Language Processing, as it requires the model to understand the nuances of language, sentiment, and style. In this paper, we propose using Large Language Models to generate Vietnamese poems from natural language prompts, thereby facilitating an intuitive process with enhanced content control. Our most efficacious model, the GPT-3 Babbage variant, achieves a custom evaluation score of 0.8, specifically tailored to the ”luc bat” genre of Vietnamese poetry. Furthermore, we also explore the idea of paraphrasing poems into normal text prompts and yield a relatively high score of 0.718 in the ”luc bat” genre. This experiment presents the potential for cross-Language poem-to-poem translation with translated poems as the inputs while concurrently maintaining complete control over the generated content.
@@ 2. Related works
Early attempts at poem generation often encountered the critical challenge of shallow content comprehension. These models relied heavily on pre-defined generation templates (inproceedings, ), term-based retrieval and reorganization to fit tonal and rhyme requirements (inproceedings2, ), or translation-based statistical approach for picking the most probable following sentence (inproceedings3, ), resulting in poems lacking creativity, coherency and controllability. Recent years have witnessed significant advancements in this field, evident in models utilizing recurrent neural networks (RNNs) like those proposed by Wang et al. (wang2016chinese, ), as well as the application of attention mechanisms (wang2016chinese2, ). These approaches have demonstrably enhanced the fluency and meaningfulness of generated poems. While Tuan Nguyen et al. (nguyen2021spgpt2, ) stands as the sole recent contribution to Vietnamese poetry generation, utilizing GPT-2 as the core foundation, they are still limited to predefined themes and starting keywords.
But eversince the impressive breakthroughs achieved by large language models like OpenAI’s ChatGPT (GPT-3.5) and GPT-3 (brown2020language, ), the field of NLP has been propelled to the forefront of public attention. Notably, ChatGPT’s remarkable skill in comprehending the intricacies and abstractions of casual language presents a compelling opportunity to revisit the domain of complex linguistic compositions such as poetry. This paper proposes a prompt-based approach that integrates the capabilities of LLMs to push the boundaries of Vietnamese poetry generation.
@@ 5.1. Experiment setup
(7) $genre=\text{Classifier}(\delta_{l_{0}},\delta_{l_{1}},...,\delta_{l_{n}})$

### 04 Jan 2024 v3

@@ Abstract.
Poetry generation has been a challenging task in the field of Natural Language Processing, as it requires the model to understand the nuances of language, sentiment, and style. In this paper, we propose using Large Language Models to generate Vietnamese poems of various genres from natural language prompts, thereby facilitating an intuitive process with enhanced content control. Our most efficacious model, the GPT-3 Babbage variant, achieves a custom evaluation score of 0.8, specifically tailored to the ”luc bat” genre of Vietnamese poetry. Furthermore, we also explore the idea of paraphrasing poems into normal text prompts and yield a relatively high score of 0.781 in the ”luc bat” genre. This experiment presents the potential for cross-Language poem-to-poem translation with translated poems as the inputs while concurrently maintaining complete control over the generated content.
@@ 2. Related works
Early attempts at poem generation often encountered the critical challenge of shallow content comprehension. These models relied heavily on pre-defined generation templates (inproceedings, ), term-based retrieval and reorganization to fit tonal and rhyme requirements (inproceedings2, ), or translation-based statistical approach for picking the most probable following sentence (inproceedings3, ), resulting in poems lacking creativity, coherency and controllability. Recent years have witnessed significant advancements in this field, evident in models utilizing recurrent neural networks (RNNs) like those proposed by Wang et al. (wang2016chinese, ), as well as the application of attention mechanisms (wang2016chinese2, ; 9207442, ). These approaches have demonstrably enhanced the fluency and meaningfulness of generated poems. While Tuan Nguyen et al. (nguyen2021spgpt2, ) stands as the sole recent contribution to Vietnamese poetry generation, utilizing GPT-2 as the core foundation, they are still limited to predefined themes and starting keywords.
But eversince the impressive breakthroughs achieved by large language models like OpenAI’s ChatGPT (GPT-3.5) and GPT-3 (brown2020language, ), the field of NLP has been propelled to the forefront of public attention. Notably, ChatGPT’s remarkable skill in comprehending the intricacies and abstractions of casual language (ye2023comprehensive, ; sawicki2023bits, ) presents a compelling opportunity to revisit the domain of complex linguistic compositions such as poetry. This paper proposes a prompt-based approach that integrates the capabilities of LLMs to push the boundaries of Vietnamese poetry generation.
@@ 5.1. Experiment setup
(7) $genre=\text{Classifier}(\delta_{l_{0}},\delta_{l_{1}},...,\delta_{l_{n-1}})$